Lewis A. Thompson

Personal Details

Address Undisclosed for publication. Telephone Undisclosed for publication.

Email lthompson@lewiz.net Date of birth Undisclosed for publication.

Nationality British WWW http://lewiz.info/

Education

• Secondary: Hymers College, Hull (September 1996 – June 2003)

A-Level: Mathematics (Grade B), Physics (Grade B), Information Communications Technology (ICT) (Grade B). AS-Level: General Studies (Grade A), Economics (Grade B). GCSE: A* in Electronic Products; As in Mathematics, Physics, Geography, German, Business Studies, English Language and English Literature; Bs in Chemistry and Biology.

• Primary: Beverley Minster C. of E. (September 1991 – July 1996)

Employment History

• The Press Association (August 2003)

Transcribing incoming news articles from different journalists and organisations. Once completed the articles are forwarded to the appropriate newspaper.

• Buttons ‘N’ Bows (July 2002 – September 2002)

Sales assistant in a needle ware, haberdashery, craft and gift ware shop.

• Fantasia (February 1998 – September 2001)

Assistant in a party hire shop.

Computer Skills

• Advanced knowledge of both Windows and UNIX-like operating environments (including Microsoft Office).

• Excellent understanding of computer networking principles, concepts and theory, including: routing, topologies and hardware.

• Knowledge of Pascal (Delphi), PHP (PHP: Hypertext Preprocessor) and BASIC programming languages.

• Proficient in Internet technologies: personally wrote a MySQL/PHP-backended website for the OHA.

Hobbies

• I am a keen kayaker, currently of grade 3* standard and look forward to taking my formal BCU (British Canoe Union) certification in the near future.

• I am an enthusiastic cyclist. I enjoy both on and off-road cycling.

References

Interested parties can apply for a list of referees.